OpenStudy (anonymous):

Evaluate. Show all work please, 9-x(-1+7x^2) for x= -2

OpenStudy (anonymous):

\[9 - 2 \times ( -1 + 7 \times 2^2) = 9 - 2 \times ( -1 + 7 \times 4) = 9 - 2 \times (27) = 9 - 54 = -45\]